The ARC coffee company sells cashews for $5.50 per pound and peanuts for $3.50 per pound. The company wants to make five pound packages of mixed nuts and sell them for $4.00 per pound. How many pounds of peanuts should be in the mix?
Answer by richwmiller(17219)

c+p=5,
5.5*c+3.5*p=20
c=5-p
5.5*(5-p)+3.5*p=20
27.5-5.5p+3.5*p=20
-2.0*p=-7.5
p=3.75
c=5-p
c=1.25 at 5.5 and p=3.75 at 3.5
check
1.25+3.75=5
5.5*c+3.5*p=20
5.5*1.25+3.5*3.75=20
6.875+13.125=20
20.0=20
ok